OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Do we have a formula to work with, or is there more information?

OpenStudy (anonymous):

its only info

OpenStudy (amistre64):

since its in feet we use -16t^2 +initial velocity(t)+initial height

OpenStudy (amistre64):

h(t) = -16t^2 +42t

myininaya (myininaya):

go amistre! you can do it!

OpenStudy (amistre64):

lol :) max height occurs when t = 42/32

OpenStudy (amistre64):

42/32 = 21/16 right? -16(21/16)^2 +42(21/16) = abt 27.56....ft

myininaya (myininaya):

21*2=42 16*2=32

myininaya (myininaya):

so yes!

OpenStudy (amistre64):

but lagrange says it cant be done; soo..... i might be wrong ;)

myininaya (myininaya):

really? does he say that?

myininaya (myininaya):

oh this lagrange not the dead one ok

OpenStudy (amistre64):

the missing info was the physics equation fro gravity

OpenStudy (anonymous):

alright alright, you got it man, nice job

myininaya (myininaya):

amistre i thought u said the dead mathematician didn't know how to reduce fractions or it cant be done lol

OpenStudy (amistre64):

there are limitations to how much can be done with an abacus ;)

myininaya (myininaya):

lol